a factory has a weak 1% acid solution and a stronger 25% acid solution. if you need to make 2000ml of 5% acid solution to perform a certain cleaning task, how many liters of each should you use?
antifreeze a is 15% alcohol. antifreeze b is 10% alcohol. how many liters of each should be mixed in order to get 20 liters of a mixture that is 18% alcohol?

Answer by Paul(988)

1.
x+25(2000-x)=5(2000)
x-25x+50000=10000
-24x=-40000
x=1666.7 (rounded)
2000-1666.7=333.3

Hence, you need to use 1666.7ml of 1% and a 333.3ml of 25%.

2.
10x+18(20)=15(x+20)
10x+360=15x+300
-5x=-60
x=12

20-12=8

Hence, about 12L of 15% solution must be added, and about 8L of 10% solution must be added.
PAUL<>